OpenStudy (anonymous):

the total area of a rectangle is 70.4375. the perimeter is 36. write the dimensions of the length and the width

OpenStudy (anonymous):

Call the short side of the rectangle x, and the longer side y. Can you think of some equations that must be satisfied. E.g. 2x+2y=36 (this is for the perimeter) What is the other one?

OpenStudy (anonymous):

L x W =70.4375 (equation1) 2L +2W=36 (equation2) Solve these by simultaneous equation:)

OpenStudy (anonymous):

x X y= 70.4375

OpenStudy (anonymous):

I tried to solve for them and ended up completing the square... which did not go very well

OpenStudy (anonymous):

So your two equations are xy=70.4375 and x+y = 18 (the same as 2x+2y=36) So you could write x as \[x = \frac{ 70.4375 }{y }\] Substituting this into x+y=18 gives us 70.4375 + y^2 = 18y y^2 - 18y + 70.4375 = 0 From here I would use the quadratic formula to get two values for y, remember this is length so we can ignore the negative answer! :)

OpenStudy (anonymous):

\[\frac{ 18+\sqrt{42.25} }{ 2 }\]

OpenStudy (anonymous):

Yeah well done, which is actually just 12.25. This is your y. Now you can work out x by subbing that back in and you have your answer!

OpenStudy (anonymous):

ok thanks